show me how to do x=3/4-4/5
Answer by lynnlo(4176)

x=3/4-4/5
SUBTRACT NUMERATOR FROM NUMERATOR==========3-4====1
MULTIPLY DENOM=============================4*5====20
X=-1/20
SOLVE LINEAR EQUATION X+1/20=0
THIS IS EQUIVALENT TO X+0.05=0
SOLUTION X=-0.05